Vs(t) (+) 0.1H Zeq If Vs(t) = 10 cos (10t - 90°) V, determine:

If Vs(t) = 10 cos (10t - 90o) V,

determine: brief description and legible formulas and variables
Step 1: The voltage phasor VSF
Step 2: The equivalent impedance Zeq
Step 3: The phasor current IF
Step 4: The steady state current I(t)ss
Step 5: The complex power (S = ½VSFIF* = P + jQ).
Step 6: The average (P) and reactive power (Q).
Step 7: The power factor and sketch of the power triangle.
Answer #1

solni- Nst) CO I) 30.1 H Given Vs(t) = 10 cos(106-90°) V | = 10 Cor Go-lot) v | Sinle Cost) Caso step 1: The voltage phasorStep 2: The equivalent impedance zeel Zeq> W R2+x? _tante R- de , L= 0,11 XL: 271f1= 2,74% 10, X0.1 25 = 12 Reqlz Ni+12 = arstep3: Phasor current IF IF: V 10 sin 10+ Z 1tj = 10 sin 100 2 L45 = 10 sin (106-450) = 552 sin (10t - 450) 5.7.04 sin (107-4L (10 sin lot), Iv [R+S4] 10X10 satice z IS) [1+0.13] I(S) z 100 (1+0.15) 84+100) z 1000 (5+10) (5271000) LITAR lim I(t) #to= { x 10 siniot.(4-04 sin(106-45-))* = {x10x4.07.ejo, e145 = 25v2 ej 45 = 2512 ( cos 45 +jsin 450) = 2502 ( 4 /3 + 2) z 25 (stepI: Power factors Cos 45= 0,709 lagging Power triangle IX, zin 10 - R=12